# Internal Deployment Gaps in AI Regulation

### v1

@@ 1. Introduction
EU AI Act + GPAI Code of Practice Research and development activities are exempt, but internal operational use may or may not be covered—the boundary is unclear; companies can also define their own acceptable risk thresholds Reassessment required when capabilities change significantly, but the company decides what counts as “significant” Detailed capability assessments only required when releasing systems externally; internally-deployed systems need only high-level policy documentation
@@
    We demonstrate that these gaps appear across the EU AIA and related GPAI COP, California’s SB 53, New York’s RAISE Act, and the US AIREA proposal, despite different jurisdictional contexts and design processes.
@@ 3.1. Scope and Coverage Gaps
R&D exemptions blur the boundary between exempt research and regulated operations. The EU AI Act presents an internal tension. Article 2(8) exempts “research, testing or development activity…prior to being placed on the market or put into service,” yet Article 3(11) defines “putting into service” to include “own use”, explicitly contemplating internal deployment. Scholars have noted that such research exemptions “do not clearly exclude commercial research and development activities,” enabling firms to “circumvent AI Act requirements by structuring work as R&D” (Yew et al., 2025). The GPAI Code of Practice attempts to close this loophole by shifting the regulatory trigger from a discrete market event to a continuous process: Recital 29 and Safety and Security Measure 1.2 require signatories to “assess and mitigate systemic risks…along the entire model lifecycle,” mandating that a Safety and Security Framework be established before market placement.
@@
Even when internal use is covered, companies may be able to unilaterally determine what counts as acceptable risk. Even where regulations nominally cover internal use, they may permit companies to define their own compliance boundaries. The GPAI Code of Practice, while extending obligations across the model lifecycle, empowers signatories to “define appropriate systemic risk tiers” and “systemic risk acceptance criteria” within their own Safety and Security Frameworks (Safety and Security Measure 4.1). By allowing companies to set thresholds for what constitutes “acceptable” internal risk, the regulation permits them to effectively determine the scope of their own operational constraints. A model deployed internally for research automation is technically covered, but can be practically exempted from strict mitigations if the developer defines internal environments as inherently low-risk within their framework.
@@
Common challenges: gaps through which internal deployment may escape coverage. Each regulation analyzed contains mechanisms through which internal deployment may escape intended coverage: broad R&D exemptions, narrow deployment definitions, self-determined risk thresholds, or combinations of all three. Even SB 53 and RAISE, which most directly address internal use, leave key boundaries undefined—most notably the threshold for “extensive” internal use that triggers regulatory obligations. The common thread is that regulations designed around a clear develop-then-deploy sequence struggle to classify systems that blur this boundary regarding performing operational work and simultaneously contributing to ongoing development.
@@ 3.2. Static Compliance Criteria for Continuously Evolving Systems
The EU framework ties reassessment to capability changes, but companies determine what counts as material changes. The EU’s GPAI Code of Practice represents the most sophisticated attempt to address continuous evolution. Rather than relying only on fixed intervals, Safety and Security Measure 1.2 mandates “lighter-touch model evaluations” along the “entire model lifecycle” whenever specific thresholds are crossed: changes in “training compute…inference compute, and/or affordances.” Safety and Security Measure 7.6 requires Model Report updates when capabilities change materially, citing “access to additional tools” or “increase in inference compute” as triggering events. This approach ties assessment to actual system changes rather than merely calendar dates.
However, the trigger mechanism introduces subjective materiality as a new gap. Re-assessment obligations activate only when providers have “reasonable grounds to believe” their previous safety justification has been “materially undermined.” This standard delegates threshold-setting to the regulated entity. Cumulative capability gains through daily fine-tuning, expanded tool access, or integration with other internal systems may never cross a self-defined materiality threshold, particularly if the provider characterizes each incremental change as immaterial in isolation. The EU framework replaces static compliance with event-driven compliance, but the events themselves remain self-reported and self-evaluated.
@@ 4.2. Approaches and Tradeoffs
Governments could create incentive structures encouraging companies to exceed minimum reporting requirements. Governments could create incentive structures encouraging companies to exceed minimum reporting requirements, sharing detailed capability evaluations, near-miss incidents, and lessons learned. Incentives might include liability protections for good-faith disclosures of security incidents or capability concerns, priority consideration for government contracts or partnerships, regulatory streamlining such as fast-track processes for companies demonstrating transparency, and access to government compute resources or specialized testing infrastructure at national laboratories. This approach leverages companies’ reputational interests and could build broader situational awareness across government without mandating disclosure of all sensitive information. This approach leverages companies’ reputational interests and could build situational awareness without mandating disclosure of all sensitive information. The tradeoff is that companies already committed to transparency will participate, while actors taking greater risks will not.
@@ Acknowledgments
We are grateful to Oscar Delaney, Ashwin Acharya, Megan Cansfield, Elias Groll, Martin Fukui, and Jake Steckler for discussion and feedback.

### v2

@@ 1. Introduction
EU AI Act + GPAI Code of Practice Provisions suggest internal deployment triggers coverage when AI systems are ”put into service” for ”own use.” The main exception is narrow: Article 2(6) exempts only systems specifically developed and deployed for the sole purpose of scientific R&D. GPAI models trigger coverage upon integration into AI systems. For GPAI models with systemic risk, reassessment required when capabilities change materially, but providers determine what constitutes a ”material” change to their safety justification Detailed Model Reports required when GPAI models are placed on the market; for internal systems, primary documentation is the high-level Safety and Security Framework. Serious Incident Reporting applies across the model lifecycle.
@@
    We analyze whether these gaps appear across the EU AIA and related GPAI COP, California’s SB 53, New York’s RAISE Act, and the US AIREA proposal, given different jurisdictional contexts and design processes.
@@ 3.1. Scope and Coverage Gaps
The EU AI Act’s scope provisions offer multiple pathways for covering internal deployment, though key boundaries remain untested. Article 3(11) defines ”putting into service” to include supply of an AI system ”for own use,” which can be interpreted to encompass internal deployment. When read alongside Articles 2(1)(a)-(c), this suggests that deploying an AI system internally in the EU triggers the Act’s obligations (Pistillo, 2025). Article 2(8), which excludes ”research, testing or development activity… prior to being placed on the market or put into service,” may not create an exception for internal deployment—rather as it describes activities that precede any deployment, whether internal or external. The narrower exception is Article 2(6), which excludes AI systems ”specifically developed and put into service for the sole purpose of scientific research and development.” However, this exemption is intended to apply to systems built for scientific R&D, such as model organisms for safety research or mechanistic interpretability (Pistillo, 2025). Using AI systems to accelerate commercial model development should not qualify, as this constitutes product-oriented research under the Act’s framework.
Safety and Security Measure 1.2 requires signatories to assess and mitigate systemic risks ”along the entire model lifecycle,” including development phases before market placement. The Code of Practice applies only to GPAI models already within the Act’s scope; a question is whether internally deployed GPAI models fall within scope as Article 2(1)(a) references GPAI models only in the context of ”placing on the market,” not ”putting into service.” But Recital 97 notes that when a provider integrates a GPAI model into an AI system that is put into service, the model ”should be considered to be placed on the market.” Since internal deployment typically requires system integration (adding affordances, tool access, and infrastructure connections), most operationally deployed internal GPAI models would trigger coverage through this pathway (Pistillo, 2025).
For GPAI models with systemic risk, implementation details remain provider-driven within the Act’s existing scope. For GPAI models with systemic risk that already fall within the Act’s scope, the Code of Practice grants signatories discretion over implementation details. Specifically, Safety and Security Measure 4.1 permits signatories to ”define appropriate systemic risk tiers” and ”systemic risk acceptance criteria” within their Safety and Security Frameworks. This self-definition operates only within boundaries already established by the Act and it does not permit providers to determine whether their models fall within scope in the first place. However, it does create latitude in how stringently providers apply mitigations to covered internal systems, potentially allowing developers to characterize internal environments as lower-risk contexts warranting less intensive safeguards.
@@
Common challenges: gaps through which internal deployment may escape coverage. Each regulation analyzed contains mechanisms through which internal deployment may escape intended coverage: broad R&D exemptions, narrow deployment definitions, self-determined risk thresholds, or combinations of all three. Even those which most directly address internal use, leave key boundaries undefined—most notably the threshold for “extensive” internal use that triggers regulatory obligations. The common thread is that regulations designed around a clear develop-then-deploy sequence struggle to classify systems that blur this boundary regarding performing operational work and simultaneously contributing to ongoing development.
@@ 3.2. Static Compliance Criteria for Continuously Evolving Systems
For GPAI models with systemic risk, the EU framework ties reassessment to capability changes, but materiality determinations remain provider-driven. For GPAI models with systemic risk, the Code of Practice moves beyond fixed assessment intervals; Safety and Security Measure 1.2 mandates evaluations ”along the entire model lifecycle” when specific thresholds are crossed: changes in training compute, inference compute, or affordances. Safety and Security Measure 7.6 requires Model Report updates when capabilities change materially, citing ”access to additional tools” or ”increase in inference compute” as triggering events. This event-driven approach ties assessment to actual capability changes rather than calendar dates, but introduces interpretive discretion. For example, reassessment obligations activate when providers have ”reasonable grounds to believe” their previous safety justification has been ”materially undermined”—a standard that relies on provider judgment. For internally deployed systems undergoing continuous fine-tuning, expanded tool access, or integration with other internal systems, cumulative capability gains may not trigger reassessment if each incremental change is characterized as immaterial in isolation. The framework addresses continuous evolution more directly than fixed-interval approaches, but verification depends on provider self-reporting rather than independent monitoring.
@@ 4.2. Approaches and Tradeoffs
Extending existing frameworks to internal deployment need not require regulatory reinvention. For AI systems already subject to obligations when deployed externally, the core argument for internal deployment oversight is straightforward: internal deployment of comparable systems should trigger the same security and safety requirements that developers already implement for external offerings. Oversight of internal deployment is not about constraining innovation or creating new compliance burdens, but about ensuring that risk management practices already deemed necessary for external systems are not bypassed simply because deployment occurs behind company walls.
@@ Acknowledgments
We are grateful to Matteo Pistillo, Oscar Delaney, Ashwin Acharya, Megan Cansfield, Elias Groll, Martin Fukui, and Jake Steckler for discussion and feedback.